find the equation of a line through point(3,1) slope =-2
Answer by duckness73(47)

There are many ways to solve this, but I think that I would use the "slope intercept" form of an equation, that is, y = mx + b. From the information provided, you know the slope m (that's -2), you know x (that's 3) and you know y (that's 1). So, you can figure out the intercept b. You have:
y = mx + b
1 = (-2)*3 + b
1 = -6 + b
7 = b
So, going back to the slope intercept form of the equation, you can make the substitutions for m and b:
y = mx + b
y = -2x + 7
Answer: y = -2x + 7